### v1

@@ 4.1 Example 1: MSH Baseline Comparison
Here we aim to replicate the results from MSH, who compared experimental measurements with simulations for a single skate skier on an unnamed 4.2 km course. The course is defined by 50 points with $(s,z)$ coordinates taken from a 2D plot of elevation versus skier distance [19, Fig. 1]. The data are easily extracted from the given figure using the online data extraction tool WebPlotDigitizer [21], by first uploading the plot image, then calibrating the coordinate axes, manually selecting data points using mouse-clicks, exporting the $(s_{i},z_{i})$ coordinates to a CSV file, and reading the data into Matlab with csvread.
@@
[…]
@@
MSH provide no indication whether their course is FIS-rated, but it is nonetheless interesting to verify whether it satisfies any of the homologation criteria in Section 2.4. Starting with criterion H2 for example, the arc length of the Hermite spline interpolant is 4218 m which sits within 0.5% of the stated 4.2 km length. The average gradient is 8.1% (computed as the mean value of $|\theta|$), which sits within the allowable range of 6–14% (H3). The point-wise gradients also lie between --23 and $+$24%, which exceed the FIS limit of $\pm$18%, but only on short sections and so this is also allowed (H4). Being able to easily test these criteria is a useful by-product of our interpolation approach.
@@
Figure 5: Simulations of the MSH baseline case. (a, Top Left) Elevation profile shown in terms of $(\xi,z)$ coordinates (black points) extracted from [19, Fig. 1], and interpolated with a Hermite (red, solid line) and linear spline (black, dotted line). (b, Top Right) Simulated speeds for skiers using the two splines, with the average speed of 5.1 m/s shown as a horizontal line. (c, Bottom Left) Forces (per unit mass) throughout the Hermite spline course, including $F_{p}$, $F_{g}$, $F_{s}$, $F_{d}$ and ${F}_{\text{\emph{tot}}}$. (d, Bottom Right) Zoomed view of the force plot on the steep uphill section with $1000\lesssim\xi\lesssim 1150$, where the skier accelerates to a roughly constant speed after which ${F}_{\text{\emph{tot}}}\approx 0$.
We are now ready to compare our model with the baseline test case from MSH. For a linear spline interpolant, the MSH simulation required a time of 815 s to complete which they found differed from the experimentally measured time by 2 seconds (although they did not state whether the actual time was 813 or 817 s). Applying our Matlab code to solve the ODE system (5)–(6) with the same linear interpolant requires 813 s of skiing time, which is within the range reported by MSH. They also provided no details about their ODE solver, so we can expect that some discrepancies will arise from differences in the numerical approximations being used. One advantage of using a robust ODE solver like Matlab’s ode15s is that it integrates seamlessly through derivative discontinuities in the ODE right hand side arising at spline points.
Although MSH performed most simulations on a piecewise linear course, they did include a single run with a cubic spline interpolant for which their skier takes roughly 20 s longer (or 835 s, estimated from [19, Fig. 16]). In comparison, when we run a second simulation with the Hermite spline interpolant, the skiing time increases to 823 s. This time lies roughly midway between those for MSH’s linear/cubic results, which is to be expected since the cubic spline should give the longest arc length (owing to its more oscillatory nature) and the linear spline the shortest.
@@
Consider this quote from Carlsson et al. [4]: “Generally a smaller skier is favored on uphills and the larger skier is favored on dowhills and the flat.” Investigate by varying skier mass $\pm$15 kg from the baseline value and comparing the finish times with those in MSH [19, Fig. 11]. Changing mass by itself ignores the fact that more muscular (heavier) skiers can generate higher peak propulsion force. So, repeat your simulations with ${P}_{\text{\emph{max}}}=275\pm 55$ W, assuming the same two skiers generate $\pm$20of the baseline power. This connection between mass and power is discussed further by Carlsson et al. [5].
@@ 4.2 Example 2: Welde’s Measurements of Elite Racers
The authors supply a 2D elevation plot [25, Fig. 2] for their course, which is divided into three laps of length 5 km each, with laps 1 and 3 being repeated runs of the same loop. The $(\xi,z)$ coordinates can be extracted directly from their course plot using WebPlotDigitizer, in contrast with the $(s,z)$ plot from Example 1. Because this course is so much longer than the previous example, we decided to exploit the “automatic extraction” feature of WebPlotDigitizer rather than manually selecting the points. This generates a CSV file with 203 data points that are used to build the Hermite spline interpolant pictured in Fig. 6a. It is important to note that the Welde elevation plot is under-resolved relative to that from Example 1, not only since the course is over three times longer, but also because the $z$-axis scale in the plot image is so much smaller. The reduced resolution introduces several anomalies in the spline interpolant that are easily identified by comparing laps 1 and 3 in Fig. 6a, and focusing on the sections labelled A,B,C.
@@
Welde’s experiment was based on a high-level race during the Norwegian national championships, but the authors did not indicate whether the course was FIS-certified; nonetheless, we will still discuss whether the homologation requirements are met. Criteria H2 and H3 are easily satisfied since the arc length of the spline interpolant is 15048 m (within 0.3% of the stated length) and the average gradient is 8.4% (similar to MSH). However, the gradient varies between --69 and $+$47% over the course, based on a fine-grid sampling of the spline $\Theta(\xi)$, which indicates some steep up- and downhills that might not pass FIS certification. On closer inspection, none of these gradient anomalies occur at or near spline data points but appear instead as spikes or oscillations near the center of a spline interval. This suggests that these spikes are an artifact of the spline interpolation, which likely arise from data extraction errors, or differentiating the spline, or both. But despite certain fine-scale course features that may not be accurately captured, they are few in number and restricted to relatively small sections, and so we are still confident that the Hermite spline provides a reasonable representation of the actual course.
Table 2: Summary of Welde’s data from 36 racers in a 15 km classic ski race in Tromsø, Norway [25, Supporting data file S1]. The rows list skiers divided into two groups of “slow” and “fast” skiers numbering 18 each, with the winner’s result listed separately. The columns list race time (in s) and average speed (in m/s) for the full course, with speeds for laps 1 and 3 included separately. The final column indicates the power parameter ${P}_{\text{\emph{max}}}$ needed in order that our model simulations match the time and speed to ski the full course.
@@
We now describe some specific race results reported by Welde and summarized in Table 2. They observed that the bottom 18 “slow” skiers completed the race in an average time of 2560 s (with average speed 5.86 m/s) whereas the top 18 “fast” skiers finished in 2381 s (average speed 6.30 m/s). The difference in performance between fast and slow skiers is largely a function of their fitness level, technique used, and more specifically the power levels they can sustain throughout a race. We can incorporate this fast/slow difference in our model by manually adjusting the parameter ${P}_{\text{\emph{max}}}$ in the power function (10) until a simulation yields the desired race time. Using this approach, we find that choosing ${P}_{\text{\emph{max}}}=392$ matches the average time for fast skiers is matched (to within 1 s), whereas the slow time is obtained with ${P}_{\text{\emph{max}}}=346$. Welde also singles out the race winner who finished well ahead of all other competitors with a time of 2268 s, that is captured (to within 1 s) by our model by taking ${P}_{\text{\emph{max}}}=434$. This clearly identifies the winner as an outlier, since the difference in maximum power between the winner and average fast skier is almost as large as the difference between the fast and slow groups. This large discrepancy is most likely attributable to the winner’s exceptional technique, based on Welde’s observation that the winner was the only skier who used double-poling (a very energy-intensive technique) throughout the entire race.
@@
The simulations for the winner are depicted in Fig. 6b in terms of the skier speed during laps 1 and 3, which exhibit noticeable variations especially in the three highlighted sections labelled A,B,C. Since these are identical course sections, these differences are most likely due to variations in the spline interpolants for laps 1 and 3. The average speed is indicated by a horizontal line in Fig. 6b, which is essentially identical at 6.7 m/s for both laps. This should be contrasted with the Welde results in Table 2, where the winner’s average speed drops by a wide margin of 5% between the first and last laps. The discrepancies identified in the speed plots for the two laps are too small and short-lasting to cause such a large reduction in speed. Instead, this slowdown between laps 1 and 3 are described by Welde as “positive pacing” and can be thought of as a consequence of muscle fatigue, in which the athlete slows down throughout a race as they exhaust their energy stores and muscles tire. As currently formulated, our model with constant ${P}_{\text{\emph{max}}}$ is not capable of capturing this fatigue effect. But we could emulate racer fatigue in a naïve way by allowing the peak power to decrease with time (or energy expenditure) throughout the simulation. A simple approach to incorporating this effect is explored in Exercise 12, but detailed experimental investigations of racer fatigue would be required in order to build a realistic model that convincingly captures Welde’s results.
Exercise 12.
@@
Notice from Table 2 that skiers are 5–6% slower on lap 3 compared to lap 1, which suggests they may be experiencing fatigue owing to their muscles tiring throughout the course. A simple way to incorporate this effect is to assume that halfway through the race (for $\xi\geqslant 7500$ m), maximum power drops from ${P}_{\text{\emph{max}}}$ to $c{P}_{\text{\emph{max}}}$, where the constant $10<c<$ 1 can be thought of as a “fatigue factor”. Modify skirun2d.m to incorporate muscle fatigue in this manner. Then adjust the values of $c$ and ${P}_{\text{\emph{max}}}$ until the total race time for the “winner” is the same as in Table 2, and their speed on lap 3 is 5–6% slower than on lap 1.
@@ 4.3 Example 3: Simulating the Toblach Ole Course in 2D
After computing the arc length and inclination angle (using the code setup3d.m), it is easy to verify that the Ole course satisfies the homologation criteria H2–H4. The length of the spline is 4157 m and the average gradient is 7.9%, both of which are well within the stated limits (for H2 and H3). The pointwise gradients also remain between $\pm$18as required by H4, except for a few short sections where it spikes to 40%. Regarding criterion H5, the last 100 m of the course are almost completely flat which is not ideal but may still be considered acceptable; however, the start begins immediately with a 5 uphill slope that clearly violates H5 (this issue was addressed previously in Exercise 3).
@@
Exercise 13.
On the Dolomiti NordicSki website [6], find the elevation profile plot for the 1.4 km “Stadium Track” course (also called “Albert Sprint”), which is a sequence of two hills having roughly similar size and shape. Use WebPlotDigitizer [21] to extract a sample of $(\xi,z)$ points with roughly 100 m spacing, write them to a CSV file, and use them to simulate a race with the baseline parameters. Next, consider a second manufactured course built from a sum of two “Gaussian bump” functions of the form $a\exp\left(-\frac{1}{c}(\xi-b)^{2}\right)$, and tune the parameters $a,b,c$ to most closely match the first course. Generate CSV data with similar resolution and repeat the simulation. How do the two results compare?
@@ 5.1 Curvature and Braking Force
In steep downhill turns, skiers reduce their speed by applying a braking force using one of two primary techniques called the skid turn and step turn [^16] In a skid turn, skis are held parallel at a slight angle to the direction of motion, and ski edges are scraped against the snow to both decelerate and turn. In a step turn, a skating motion is combined with rapid side-to-side stepping to maintain speed as much as possible during the turn and dissipate less energy. There is a third technique called the snowplow, which is an extreme version of the skid turn and is more of a beginner’s strategy that is seldom applied by experienced racers. An in-depth discussion of turning techniques can be found in the Nordic Ski Lab article “Nordic downhill ski techniques for safety and speed”.. In both cases, the braking turn generates a force opposing forward motion, which we assume following Sandbakk et al. [22] is proportional to centripetal acceleration so that $F_{b}=\gamma a_{c}=\gamma\kappa v^{2}$, where $\gamma$ is a dimensionless braking coefficient.
It is unrealistic for an athlete to apply a braking force throughout an entire race, even if the force is mostly very small. Therefore, $F_{b}$ should be “thresholded” so that it is zero except within those downhill sections where centripetal acceleration exceeds some braking threshold ${a}_{\text{\emph{c,min}}}$:
@@
Before moving on, we should mention that some Nordic ski models [5, 20, 24] incorporate an extra contribution to snow friction arising from the component of the centripetal force acting within the $\xi,z$-plane along the direction of travel [^18] Vector calculus tell us that the (signed) curvature for a 2D space curve $z(\xi)$ is given by $\kappa_{{}_{2D}}(\xi)=z^{\prime\prime}(\xi)/(1+z^{\prime}(\xi))^{3/2}$. Note that this formula is equivalent to Eq. (13), except that it is a signed quantity that is positive/negative depending on whether the 2D ’ curve is concave up/down., which is always directed normal to the snow surface. MSH observed that they “find the greatest differences between the simulation and the experimental data on terrain with rapidly changing curvature” [19], which we also see at isolated instants in time. However, because most ski courses have a balance between up/down-hill sections, the +/+/-+ / - contributions from this curvature term tend to cancel out on average, leading to a negligible net effect on the overall skiing time. Therefore, we have chosen to ignore the effects of this vertical curvature on snow friction (which is also consistent with the MSH model [19]).

### v2

@@ 4.1 Example 1: MSH Baseline Comparison
Here we aim to replicate the results from MSH, who compared experimental measurements with simulations for a single skate skier on an unnamed 4.2 km course. The course is defined by 50 points with $(s,z)$ coordinates taken from a 2D plot of elevation versus skier distance [24, Fig. 1]. The data are easily obtained from the given figure using the online data extraction tool WebPlotDigitizer [27], by first uploading the plot image, then calibrating the coordinate axes, manually selecting data points using mouse-clicks, exporting the $(s_{i},z_{i})$ coordinates to a CSV file, and reading the data into MATLAB with csvread. The interested reader is welcome to investigate other approaches for obtaining course data using a package such as GRABIT (distributed through the MATLAB File Exchange [7]) or the built-in command ginput, which would then permit both the data extraction and simulation steps to be performed within the MATLAB environment.
@@
$\xi_{i}=\xi(\alpha_{i})\approx\int_{0}^{\alpha_{i}}\displaystyle\sqrt{S_{\alpha}^{2}-Z_{\alpha}^{2}}\;d\alpha,$
@@
MSH provide no indication whether their course is FIS-rated, but it is nonetheless interesting to verify whether it satisfies any of the homologation criteria in Section 2.4. Starting with criterion H2 for example, the arc length of the Hermite spline interpolant is 4218 m which sits within 0.5% of the stated 4.2 km length. The average gradient is 8.1% (computed as the mean value of $|\theta|$), which sits within the allowable range of 6–14% (H3). The point-wise gradients also lie between -23 and ++24%, which exceed the FIS limit of $\pm$18%, but only on short sections and so this is also allowed (H4). Being able to easily test these criteria is a useful by-product of our interpolation approach.
@@
Figure 5: Simulations of the MSH baseline case. (a, Top Left) Elevation profile shown in terms of $(\xi,z)$ coordinates (black points) extracted from [24, Fig. 1], and interpolated with a Hermite (red, solid line) and linear spline (black, dotted line). (b, Top Right) Simulated speeds for skiers using the two splines, with the average speed of 5.1 m/s shown as a horizontal line. (c, Bottom Left) Forces (per unit mass) throughout the Hermite spline course, including $F_{p}$, $F_{g}$, $F_{s}$, $F_{d}$ and ${F}_{\text{\emph{tot}}}$. (d, Bottom Right) Zoomed view of the force plot on the steep uphill section with $1000\lesssim\xi\lesssim 1150$, where the skier accelerates to a roughly constant speed after which ${F}_{\text{\emph{tot}}}\approx 0$.
We are now ready to compare our model with the baseline test case from MSH. For a linear spline interpolant, the MSH simulation required a time of 815 s to complete which they found differed from the experimentally measured time by 2 seconds (although they did not state whether the actual time was 813 or 817 s). Applying our MATLAB code to solve the ODE system (5)–(6) with the same linear interpolant requires 813 s of skiing time, which is within the range reported by MSH. They also provided no details about their ODE solver, so we can expect that some discrepancies will arise from differences in the numerical approximations being used. One advantage of using a robust ODE solver like MATLAB’s ode113 is that it integrates seamlessly through derivative discontinuities in the ODE right-hand side arising at spline points.
Although MSH performed most simulations on a piecewise linear course, they did include a single run with a cubic spline interpolant for which their skier takes roughly 20 s longer (or 835 s, estimated from [24, Fig. 16]). In comparison, when we run a second simulation with the Hermite spline interpolant, the skiing time increases to 823 s. This time lies roughly midway between those for MSH’s linear/cubic results, which is to be expected since the cubic spline should give the longest arc length (owing to its more oscillatory nature) and the linear spline the shortest.
@@
Consider this quote from Carlsson et al. [4]: “Generally a smaller skier is favored on uphills and the larger skier is favored on dowhills and the flat.” Investigate by varying skier mass $\pm$15 kg from the baseline value and comparing the finish times with those in MSH [24, Fig. 11]. Changing mass by itself ignores the fact that more muscular (heavier) skiers can generate higher peak propulsion force. So, repeat your simulations with ${P}_{\text{\emph{max}}}=275\pm 55$ W, assuming the same two skiers generate $\pm$20% of the baseline power. This connection between mass and power is discussed further by Carlsson et al. [5].
Exercise 12.
Compare the accuracy and computational cost when MATLAB’s ODE solvers ode113, ode45, ode15s and ode23s are applied to the “baseline” MSH example. Use the difference in race times as a simple measure of accuracy, and compare computational cost using the built-in functions tic, toc. Discuss the results, paying particular attention to the algorithms implemented in the four solvers as described in the MATLAB documentation.
@@ 4.2 Example 2: Welde’s Measurements of Elite Racers
The authors supply a 2D elevation plot [34, Fig. 2] for their course, which is divided into three laps of length 5 km each, with laps 1 and 3 being repeated runs of the same loop. The $(\xi,z)$ coordinates can be extracted directly from their course plot using WebPlotDigitizer, in contrast with the $(s,z)$ plot from Example 1. Because this course is so much longer than the previous example, we decided to exploit the “automatic extraction” feature of WebPlotDigitizer rather than manually selecting the points. This generates a CSV file with 203 data points that are used to build the Hermite spline interpolant pictured in Fig. 6a. It is important to note that the Welde elevation plot is under-resolved relative to that from Example 1, not only since the course is over three times longer, but also because the $z$-axis scale in the plot image is so much smaller. The reduced resolution introduces several anomalies in the spline interpolant that are easily identified by comparing laps 1 and 3 in Fig. 6a, and focusing on the sections labelled A,B,C.
@@
Welde’s experiment was based on a high-level race during the Norwegian national championships, but the authors did not indicate whether the course was FIS-certified; nonetheless, we will still discuss whether the homologation requirements are met. Criteria H2 and H3 are easily satisfied since the arc length of the spline interpolant is 15048 m (within 0.3% of the stated length) and the average gradient is 8.4% (similar to MSH). However, the gradient varies between -69 and ++47% over the course, based on a fine-grid sampling of the spline $\Theta(\xi)$, which indicates some steep up- and downhills that might not pass FIS certification. On closer inspection, none of these gradient anomalies occur at or near spline data points but appear instead as spikes or oscillations near the center of a spline interval. This suggests that these spikes are an artifact of the spline interpolation, which likely arise from data extraction errors, or differentiating the spline, or both. But despite certain fine-scale course features that may not be accurately captured, they are few in number and restricted to relatively small sections, and so we are still confident that the Hermite spline provides a reasonable representation of the actual course.
Table 2: Summary of Welde’s data from 36 racers in a 15 km classic ski race in Tromsø, Norway [34, Supporting data file S1]. The rows list skiers divided into two groups of “slow” and “fast” skiers numbering 18 each, with the winner’s result listed separately. The columns list race time (in s) and average speed (in m/s) for the full course, with speeds for laps 1 and 3 included separately. The final column indicates the power parameter ${P}_{\text{\emph{max}}}$ needed in order that our model simulations match the time and speed to ski the full course.
@@
We now describe some specific race results reported by Welde and summarized in Table 2. They observed that the bottom 18 “slow” skiers completed the race in an average time of 2560 s (with average speed 5.86 m/s) whereas the top 18 “fast” skiers finished in 2381 s (average speed 6.30 m/s). The difference in performance between fast and slow skiers is largely a function of their fitness level, technique used, and more specifically the power levels they can sustain throughout a race. We can incorporate this fast/slow difference in our model by manually adjusting the parameter ${P}_{\text{\emph{max}}}$ in the power function (10) until a simulation yields the desired race time. Using this approach, we find that choosing ${P}_{\text{\emph{max}}}=392$ matches the average time for fast skiers (to within 1 s), whereas the slow time is obtained with ${P}_{\text{\emph{max}}}=346$. Welde also singles out the race winner who finished well ahead of all other competitors with a time of 2268 s, that is captured (to within 1 s) by our model by taking ${P}_{\text{\emph{max}}}=434$. This clearly identifies the winner as an outlier, since the difference in maximum power between the winner and average fast skier is almost as large as the difference between the fast and slow groups. This large discrepancy is most likely attributable to the winner’s exceptional technique, based on Welde’s observation that the winner was the only skier who used double-poling (a very energy- intensive technique) throughout the entire race. We also note that the elite skier from Example 1 had a much lower value of ${P}_{\text{\emph{max}}}$ than all three groups in Welde’s study, which suggests that the self-chosen pace of the MSH athlete was significantly slower than their race pace.
@@
The simulations for the winner are depicted in Fig. 6b in terms of the skier speed during laps 1 and 3, which exhibit noticeable variations especially in the three highlighted sections labelled A,B,C. Since these are identical course sections, these differences are most likely due to variations in the spline interpolants for laps 1 and 3. The average speed is indicated by a horizontal line in Fig. 6b, which is essentially identical at 6.7 m/s for both laps. This should be contrasted with the Welde results in Table 2, where the winner’s average speed drops by a wide margin of 5% between the first and last laps. The discrepancies identified in the speed plots for the two laps are too small and short- lasting to cause such a large reduction in speed. This slowdown between laps 1 and 3 is described by Welde as “positive pacing” and can be thought of as a consequence of muscle fatigue, in which the athlete slows down throughout a race as they exhaust their energy reserves and muscles tire. As currently formulated, our model with constant ${P}_{\text{\emph{max}}}$ is not capable of capturing this fatigue effect; however, we could easily emulate racer fatigue in a naïve way by allowing the peak power to decrease with time throughout a simulation. This simple approach to modelling fatigue is explored in Exercise 13, although fully capturing Welde’s results would require a much deeper study that incorporates experimental observations on positive pacing (e.g., [15, 31]).
Exercise 13.
@@
Notice from Table 2 that skiers are 5–6% slower on lap 3 compared to lap 1, which suggests they may be experiencing fatigue owing to their muscles tiring throughout the course. A simple way to incorporate this effect is to assume that halfway through the race (for $\xi\geqslant 7500$ m), maximum power drops from ${P}_{\text{\emph{max}}}$ to $c{P}_{\text{\emph{max}}}$, where the constant $0<c<1$ can be thought of as a “fatigue factor”. Modify skirun2d.m to incorporate muscle fatigue in this manner. Then adjust the values of $c$ and ${P}_{\text{\emph{max}}}$ until the total race time for the “winner” is the same as in Table 2, and their speed on lap 3 is 5–6% slower than on lap 1.
@@ 4.3 Example 3: Simulating the Toblach Ole Course in 2D
After computing the arc length and inclination angle (using the code setup3d.m), it is easy to verify that the Ole course satisfies the homologation criteria H2–H4. The length of the spline is 4157 m and the average gradient is 7.9%, both of which are well within the stated limits (for H2 and H3). The pointwise gradients also remain between $\pm$18% as required by H4, except for a few short sections where it spikes to 40%. Regarding criterion H5, the last 100 m of the course are almost completely flat which is not ideal but may still be considered acceptable; however, the start begins immediately with a 5 uphill slope that clearly violates H5 (this issue was addressed previously in Exercise 3).
@@
On the Dolomiti NordicSki website [8], find the elevation profile plot for the 1.4 km “Albert Sprint” course (also called “Stadium Track”), which is a sequence of two hills having roughly similar size and shape. Use WebPlotDigitizer [27] to extract a sample of $(\xi,z)$ points with roughly 100 m spacing, write them to a CSV file, and use them to simulate a race with the baseline parameters. Next, consider a second manufactured course built from a sum of two “Gaussian bump” functions of the form $a\exp\left(-\frac{1}{c}(\xi-b)^{2}\right)$, and tune the parameters $a,b,c$ to most closely match the first course. Generate CSV data with similar resolution and repeat the simulation. How do the two results compare?
Exercise 15.
@@ 5.1 Curvature and Braking Force
In steep downhill turns, skiers reduce their speed by applying a braking force using one of two primary techniques called the skid turn and step turn [^18] In a skid turn, skis are held parallel at a slight angle to the direction of motion, and ski edges are scraped against the snow to both decelerate and turn. In a step turn, a skating motion is combined with rapid side-to-side stepping to maintain speed as much as possible during the turn and dissipate less energy. There is a third technique called the snowplow, which is an extreme version of the skid turn and is more of a beginner’s strategy that is seldom applied by experienced racers. An in-depth discussion of turning techniques can be found in the Nordic Ski Lab article “Nordic downhill ski techniques for safety and speed”.. In both cases, the braking turn generates a force opposing forward motion, which we assume following Sandbakk et al. [28] is proportional to centripetal acceleration so that $F_{b}=\gamma a_{c}=\gamma\kappa v^{2}$, where $\gamma$ is a dimensionless braking coefficient [^19] A similar assumption that braking force is proportional to centripetal acceleration is applied to a dynamic model for road cycling in [25].. It is unrealistic for an athlete to apply a braking force throughout an entire race, even if the force is mostly very small. Therefore, $F_{b}$ should be “thresholded” so that it is zero except within those downhill sections where centripetal acceleration exceeds some braking threshold ${a}_{\text{\emph{c,min}}}$:
@@
Before moving on, we should mention that some Nordic ski models [5, 26, 33] incorporate an extra contribution to snow friction arising from the component of the centripetal force acting within the $\xi,z$-plane along the direction of travel [^21] Vector calculus tell us that the (signed) curvature for a 2D space curve $z(\xi)$ is given by $\kappa_{{}_{2D}}(\xi)=z^{\prime\prime}(\xi)/(1+z^{\prime}(\xi))^{3/2}$. Note that this formula is equivalent to Eq. (13), except that it is a signed quantity that is positive/negative depending on whether the 2D ’ curve is concave up/down., which is always directed normal to the snow surface. MSH observed that they “find the greatest differences between the simulation and the experimental data on terrain with rapidly changing curvature” [24], which we also see at isolated instants in time. However, because most ski courses have a balance between up/down-hill sections, the +/+/- contributions from this curvature term tend to cancel out on average, leading to a negligible net effect on the overall skiing time. Therefore, we have chosen to ignore the effects of this vertical curvature on snow friction (which is also consistent with the MSH model [24]).